OpenStudy (anonymous):

solve the equation. 7/2x+3=x+11/2

OpenStudy (anonymous):

\[\frac{7}{2x + 3} = \frac{x +11}{2}\] Right?

OpenStudy (anonymous):

Okay, Now lets do this together.

OpenStudy (anonymous):

x doesnt go on top(on side)

OpenStudy (anonymous):

\[\frac{7}{2x +3} = x + \frac{11}{2}\] Is it correct now ?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

So 7 / 2x +3 = 2x +11/ 2 Can u cross multiply?

OpenStudy (anonymous):

14=(2x+11)(2x+3)=4x^2+28x+33 4x^2+28x+19=0 x=\[[-28\pm \sqrt{28^{2}-4*4*19}\ ]/2*4\]